OpenStudy (griffindore):

If (7^2)^p = 7^6, what is the value of p? 2 3 4 8

OpenStudy (griffindore):

c right?

OpenStudy (griffindore):

oh really ok thanks

TheSmartOne (thesmartone):

That is wrong

TheSmartOne (thesmartone):

Hint: \(\sf\Huge a^{{(b)}^{(c)}} = a^{bc}\) \(\sf\Huge x^y = x^z \to y=z\)

OpenStudy (griffindore):

is it 4?

TheSmartOne (thesmartone):

nope